OpenStudy (s):

Find y' where y=x^(2x) *x^(x^2+2) ?

OpenStudy (jhannybean):

\[f(x) = x^{2x}\cdot x^{x^2+2} = x^{x^2+2x+2}\]

OpenStudy (jhannybean):

Factor out the top first.

OpenStudy (jhannybean):

OR NO

OpenStudy (dan815):

log and implicit

OpenStudy (jhannybean):

You're just trying to find y'. Then just -.... Thanks dan -.-

OpenStudy (dan815):

i think

OpenStudy (jhannybean):

Ther is no other way, you have a base x raised to the power of the same variable. I would think so

OpenStudy (s):

yes, I am stuck with this problem for a while. I thought of some sort of substitution, but it didn't help

OpenStudy (anonymous):

Continuing\[ \ln(y) = (x^2+2x+2)\ln(x) \]

OpenStudy (anonymous):

\[ \frac{d}{dx}\ln(y) =\frac{d}{dy}\ln(y) \frac{dy}{dx}= \frac{1}{y}y' \]

OpenStudy (anonymous):

\[ \frac{y'}{y} = (2x+2)\ln(x) + (x^2+2x+2)\frac{1}{x} \]

OpenStudy (anonymous):

We can substitute \(y\) for it's function of \(x\).\[ \frac{y'}{y}=\frac{y'}{x^{x^2+2x+2}} \]

OpenStudy (anonymous):

This is how you do things with logarithmic differentiation.

OpenStudy (anonymous):

Another way to do this, is to use: \[ a^x = e^{x\ln a } \implies x^{x^2+2x+2} = e^{(x^2+2x+2)\ln(x)} \]

OpenStudy (anonymous):

Both result in: \[ y'=\left(x^{x^2+2x+2}\right)\left((2x+2)\ln(x)+\frac{x^2+2x+2}{x}\right) \]

OpenStudy (s):

Thank you very much again ! Now since you explained it all it is not as bad as it was..

OpenStudy (s):

I appreciate it !

OpenStudy (anonymous):

So I didn't do anything confusing?